OpenStudy (anonymous):

I have a Jacobi matrix (An) with all a's down the main diagonal, all b's in the diagonal above the a's and all c's in the diagonal below, everything else is zeroes...need help proving that detAn=a*detA(n-1)-bc*detA(n-2) for all n

OpenStudy (anonymous):

If you help me out with what a Jacobi matrix is, I bet we can figure this out.

OpenStudy (anonymous):

well it's actually not a jacobi, it's modified, jacobi means that whenever absolute(i-j) is bigger or equal to 2, Aij=0

OpenStudy (anonymous):

Ok, but that is essentially taken care of in the rest of the statement.

OpenStudy (anonymous):

right, what I wrote initially represents the matrix pretty well

OpenStudy (anonymous):

Have you tried expanding the determinant by minors?

OpenStudy (anonymous):

no I haven't yet

OpenStudy (anonymous):

I see how to get the first term this way. I'm having some trouble seeing the second term without writing it out, but I can see how would most likely fall out.

OpenStudy (anonymous):

I've been able to get the equation from solving detH3, but I cannot extend the equation for all n above 3

OpenStudy (anonymous):

if this is undoable, my next question asks me to find H6..

OpenStudy (anonymous):

sorry A6

OpenStudy (anonymous):

Well, if we expand by minors in the first row, then we get \[\det(A_{n})=a_{11}\left|\begin{array}{cccc}a_{22}&a_{23}&\ldots&a_{2n}\\a_{32}&a_{33}&\ldots&a_{3n}\\ \vdots&\vdots&\ddots&\vdots\\ a_{n2}&a_{n3}&\ldots&a_{nn}\end{array}\right|-a_{12}\left|\begin{array}{cccc}a_{21}&a_{23}&\ldots&a_{2n}\\a_{31}&a_{33}&\ldots&a_{3n}\\ \vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n3}&\ldots&a_{nn}\end{array}\right|\]

OpenStudy (anonymous):

\[+\ldots\pm a_{1n}\left|\begin{array}{cccc}a_{21}&a_{22}&\ldots&a_{2(n-1)}\\a_{31}&a_{32}&\ldots&a_{3(n-1)}\\ \vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&\ldots&a_{n(n-1)}\end{array}\right|\]

OpenStudy (anonymous):

Notice that lopping off the first row and first column just gives us \(A_{n-1}\), so the first term will be the \(a\det(A_{n-1})\)

OpenStudy (anonymous):

right, and lopping b gives [c b 0], [0 a b], [0 c a]

OpenStudy (anonymous):

Is that for the second one?

OpenStudy (anonymous):

yeah, sorry my comp is not running too well

OpenStudy (anonymous):

how do you make matrices?

OpenStudy (anonymous):

that's not what I got for the matrix in the second term

OpenStudy (anonymous):

sorry, that was my matrix for the second term in row 1

OpenStudy (anonymous):

row 2, column 1 is b right? and row 2 column 3 is c?

OpenStudy (anonymous):

[a b 0 0], [c a b 0], [0 c a b], [0 0 c a]

OpenStudy (anonymous):

Oh, I had the b and c switched. Okay, we're good. You're right about the second term.

OpenStudy (anonymous):

Ok, now we are going to expand the second term determinant by minors again.

OpenStudy (anonymous):

the other property that this matrix has is that Hii=a, Hi, i+1=b Hi, i-1=c

OpenStudy (anonymous):

so maybe the matrix is different for A5? not sure, they give us A4, so I don't exactly know what a5 looks like

OpenStudy (anonymous):

\[a_{12}\left|\begin{array}{cccc}a_{21}&a_{23}&\ldots&a_{2n}\\a_{31}&a_{33}&\ldots&a_{3n}\\ \vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n3}&\ldots&a_{nn}\end{array}\right|=b\left|\begin{array}{cccc}c&b&0&\ldots&0\\0&a&b&\ldots&0\\0&c&a&\ldots&0\\\vdots&\vdots&\vdots&\ddots&\vdots\\0&0&0&\ldots&a\end{array}\right|\]

OpenStudy (anonymous):

Does that look good?

OpenStudy (anonymous):

That property just says what you already said.

OpenStudy (anonymous):

yup, I keep forgetting that we have to deal with An

OpenStudy (anonymous):

Okay, so what do the rest of the terms after term 2 look like?

OpenStudy (anonymous):

they should all be zero because abs(i-j) is bigger or equal to 2, in which case the term is zero...first row is always [a b 0 0 0 0 0...]

OpenStudy (anonymous):

Right. So, now all we have to do is show that the second term is what we want it to be. Any ideas on how to do that?

OpenStudy (anonymous):

second term is always b because the property says that Hi, i+1=b, which will always be for all n, the 2nd determinant is what I dont know how to express as H(n-2)

OpenStudy (anonymous):

Right. What did we do to find the original determinant?

OpenStudy (anonymous):

minor a12

OpenStudy (anonymous):

Well, yeah, we expanded along the first row. We are going to expand by minors again. Which row or column are we going to use?

OpenStudy (anonymous):

one sec, I'm still absorbing how the first determinant is A(n-1) I think i got it

OpenStudy (anonymous):

Oh, ok, sure.

OpenStudy (anonymous):

for the 2nd determinant the first 3 entries are [c a b 0 0 0 0 0], [0 c a b 0000], [0 0 c a b 00]

OpenStudy (anonymous):

But the determinant isn't \(A_{n-a}\), it's \(\det(A_{n-1})\)

OpenStudy (anonymous):

ur right I copied it wrong

OpenStudy (anonymous):

so it's [c b 000] [0 c b 000] [ 0 0 c b 000]

OpenStudy (anonymous):

But you skip the second column because that is the column that \(a_{12}\) is in.

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

what can we take by that? I get that there is c's down a diagonal and b's down a diagonal

OpenStudy (anonymous):

Wait, no, it's what I have up there, you only lose the first \(a\). After that you start getting \(a\)'s again.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

c b 0000, 0 c a b 00000, 0 0 c a b 0000?

OpenStudy (anonymous):

I think the second on is 0ab0...

OpenStudy (anonymous):

one

OpenStudy (anonymous):

yup, ive been doing math al day haha, not working at full potential, good catch

OpenStudy (anonymous):

I know how that goes. I usually top out after an hour at a time. Haha.

OpenStudy (anonymous):

So take a look at \[\left[\begin{array}{ccccc}c&b&0&0&\ldots\\0&a&b&0&\ldots\\0&c&a&b&\ldots\\0&0&c&a&\ldots\\\vdots&\vdots&\vdots&\vdots&\ldots\end{array}\right]\] and compare it to the original matrix.

OpenStudy (anonymous):

we have to take out the c somehow, so -bc*detA(n-2), know how we could take out the c?

OpenStudy (anonymous):

Right, we take out the c by expanding along a row or column with the c in it.

OpenStudy (anonymous):

well every row has a c except the second one

OpenStudy (anonymous):

sorry how do you expand? is that taking out a row and column and then multiplying the determinant by the entry?

OpenStudy (anonymous):

mathworld.wolfram.com/Determinant.html

OpenStudy (anonymous):

if we take out the a11 we have Hn again don't we?

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

But right now we are smaller by 1 row and 1 column. After we take out another, we are smaller by 2, which is?

OpenStudy (anonymous):

it's A(n-2)...so let's add it all up....det(An)= a11*det(Hn-1)-bc(det(Hn-1)) that's what i got so far

OpenStudy (anonymous):

that last part is wrong...I don't understand how we went from Hn-1 to Hn-2 in the 2nd determinant

OpenStudy (anonymous):

sorry i keep putting H, H or A same thing

OpenStudy (anonymous):

We expanded by minors, which means we were working with a smaller matrix.

OpenStudy (anonymous):

Yeah, I get it.

OpenStudy (anonymous):

To find that second determinant, we are going to expand again.

OpenStudy (anonymous):

oh ok, but we couldn't have said -b(detHn-1) hey?

OpenStudy (anonymous):

But it's not \(\det(A_{n-1})\), it was only like that the first time because of how the structure of the matrix and what happened after we remove the *first* row and *first* column.

OpenStudy (anonymous):

ok, I get it now, we expanded and ended up with Hn-2, cool

OpenStudy (anonymous):

Right. But it's important which row or column we expand along.

OpenStudy (anonymous):

for H6 I'm guessing it's just long but straightforward

OpenStudy (anonymous):

...oh ok, thx a lot you've been a really big help

OpenStudy (anonymous):

So for the second determinant, we expand down the first column. That way, we multiply the b by c and \(A_{n-2}\). The rest of the terms b gets multiplied by are 0.

OpenStudy (anonymous):

cool, that gives our equation

OpenStudy (anonymous):

thx, now on to a Calc assignment, going to be up all night haha

OpenStudy (anonymous):

Ha, fun.

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

thanks again have a good one!

OpenStudy (anonymous):

You too.